Consider the following table, which gives a security analyst's expected return on two stocks for two particular market returns:

Market Return Aggressive Stock Defensive Stock
6 % 2.3 % 4.1 %
16 25 14


a. What are the betas of the two stocks? (Round your answers to 2 decimal places.)




b. What is the expected rate of return on each stock if the market return is equally likely to be 6% or 16%? (Round your answers to 2 decimal places.)



c. If the T-bill rate is 8%, and the market return is equally likely to be 6% or 16%, what are the alphas of the two stocks? (Negative values should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to 2 decimal places.)
Answer #1

a) Beta (Aggressive) = (25 - 2.3) / (16 - 6) = 2.27

Beta (Defensive) = (14 - 4.1) / (16 - 6) = 0.99

b) Expected Return (Aggressive) = (25 + 2.3) / 2 = 13.65%

Expected Return (Defensive) = (14 + 4.1) / 2 = 9.05%

c) Alpha is the return of market returns. Market Returns = (6% + 16%) / 2 = 11%

Alpha (Agg) = 13.65% - 11% = 2.65%

Alpha (Def) = -1.95%